After the GMOS-S detector troubleshooting earlier this semester, the physical location of the array has shifted by a small amount in the -direction. While this did not have harmful consequences, the Y-shift has the effect of the MOS mask slits as defined on the MDF appearing with a small (~20 pix) yet systematic offset with respect to the actual location of the spectra.
It has recently been noticed that the sensitivity of GMOS-N B600 observations at wavelengths blueward of 500 nm is significantly lower than predicted by the Exposure Time Calculator. Historic data of spectrophotometric standard stars show that this reduction in sensitivity is largely attributed to a decrease of the B600 grating response, which started to develop around 2013 to 2014.
IGRINS was removed today from telescope to leave way for GHOST installation. After that, IGRINS will go through a scheduled maintenance by a team from KASI, and it will back at the telescope on October 17th.
Following successful checks on Friday after the yearly engineering shutdown, IGRINS is back doing observations. The instrument will remain on the telescope until Sept 7th.
The repair attempted during July was not successful, and the problem on CCD2 (amplifier #5) persists. On the other hand, the noise patterns on CCD2 and CCD3 are gone.
The situation will remain like this during semester 22B since the only possible solution requires a major intervention, which is not planned for the current semester.
After the heavy snowtorms that affected Pachon, telescope operations have resumed, and IGRINS will be available for a short time starting tonight until the beginning of the yearly engineering shutdown.
Heavy snowfall at Cerro Pachon has completely wiped out the Zorro run originally scheduled between July 11 and 17th. We expect the telescope can be operating again on Monday 18th, and the Zorro run has been extended until Friday 22nd.
GNIRS has been reinstalled on the telescope and is ready for science use as of July 4th, 2022. All issues with the filter wheels have been fixed and other refurbishment projects have been completed. All slit spectral modes are now available in addition to imaging. The Low Resolution (natural seeing) IFU is installed and is undergoing commissioning. It will be released for science use at a later date. Fabrication of the High Resolution (AO-assisted) IFU is ongoing and is expected to arrive in Hilo at the end of 2022.
General maintenance has been scheduled for IGRINS between September 5th and 16th, when the instrument will be unavailable. A visiting team from KASI at Pachon will execute several tasks, including the replacement of one of the cold heads.
After the completion of the GHOST commissioning, the GHOST Cassegrain Unit was removed from the telescope and IGRINS was installed back on the same port last Monday. Back on sky checks were performed successfully on Tuesday night and the instrument was used immediately for science operations.